# Evaluating the Limits of QAOA Parameter Transfer at High-Rounds on Sparse Ising Models With Geometrically Local Cubic Terms

### v1

@@ I Introduction
The Quantum Approximate Optimization Algorithm (QAOA) [1, 2, 3, 4, 5] is designed to sample good solutions of combinatorial optimization problems. QAOA is relatively NISQ (Noisy Intermediate-Scale Quantum) [6] computer friendly in that it is feasible to implement the algorithm at the scale of near term noisy quantum processors (short depth, small qubit counts). There have been a considerable number of studies on the properties of QAOA using both numerical simulations and current quantum processors, in particular in comparison to existing classical algorithms that solve combinatorial optimization problems [7, 8, 9, 10, 11, 12, 13, 14]. One of the central challenges with the algorithmic study of QAOA is that scaling sampling analysis, both with respect to number of variables and the number of QAOA steps used, requires either somewhat small scale numerical simulations (limited to tens of qubits for exact statevector computations, or up to hundreds of qubits using tensor network methods or similar types of approximate numerical techniques) or quantum hardware simulations which are subject to a variety of sources of error. QAOA requires hyper-parameter selection that in turn generally requires significant compute time, especially if this process is performed on the quantum device. However, the QAOA parameter learning is distinct from many other types of variational quantum algorithms (VQAs) [15] because the number of trainable parameters does not scale, per layer, with the number of variables in the problem, meaning that parameter consistency among problems with varying problem sizes, or coefficient distributions, can be examined. It turns out that QAOA parameters do indeed exhibit parameter consistency for problem instances that share some similarity, such as similar connection structure. This property has been studied and utilized for efficient QAOA parameter finding in numerous studies  [^1] The QAOA parameter similarity property comes by several different terms, including parameter transfer, problem instance independent QAOA parameters, fixed angles, parameter concentration, parameter fixing, size invariant QAOA parameters, fixed control parameters, and transfer learning [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Many of these types of parameter transfer studies are able to condense QAOA angle usage into fixed angle formulas, at least for average cases of certain problem types and typically at fairly low depth [33]. If QAOA can be used in a non-variational learning manner, then the prospects of using QAOA for meaningfully sized problem instances on large scale error-corrected quantum computers are good – the variational learning overhead of QAOA can be quite significant. Parameter setting heuristics, namely from parameter concentration, will enable this approach to use QAOA at much larger scale [34]. There do exist other non-variational QAOA approaches, such as using linear-ramp adiabatic schedules [35, 36], which at sufficiently large $p$ guarantees optimality.
@@
Here we show through extensive numerical simulation some of the favorable characteristics of parameter transfer, and some of the limitations. More precisely, we use QAOA parameters trained on small, statevector simulation-addressable problem instances to execute full scale QAOA runs on several IBM quantum processors. Finally, we validate the QAOA angle performance using extensive tensor network simulations, in particular PEPS (Projected Entangled Pair States). The QAOA circuits and problem optimization problems that we use in this study are carefully designed to be extremely hardware compatible to the heavy-hex graph and are therefore very NISQ-friendly, see [37, 38].
@@
In this study, the scaling of QAOA is characterized using an ensemble of learned QAOA angles on small (e.g., less than $30$ qubits) problem instances whose parameters are then transferred to be used on much larger problem instances (up to $156$ qubits). The central question that is examined is how far parameter transfer QAOA angles can be used, with respect to $p$. This is an important question because although parameter transfer seems to be incredibly successful for finding good performing QAOA angles, there are limits to direct parameter training for parameter transfer – namely, for the task of transferring parameters up to larger problem sizes, better angles cannot be learned past an approximation ratio of $1$, and higher QAOA depths are generally required as problem sizes increase. It is not clear how well parameter transfer can be applied to various combinatorial optimization problem instances in general, but it is crucial to understand where transferred QAOA angles fail to perform, or perhaps work well in certain cases. Remarkably one of the findings of this study is that the intensely optimized QAOA angles, using the classical numerical statevector simulation tool JuliQAOA, recover the smooth-schedule properties of discretized quantum annealing (e.g., adiabatic)-type protocols.
@@ II Methods
The QAOA angle finding in this study is carried out using the Julia [42] based software package JuliQAOA [43], where the angle finding is performed using basin hopping optimization, which in turn uses gradient descent, and all angle optimization begins with uniformly random guesses at $p=1$. JuliQAOA is an efficient numerical simulation angle finding tool that can heuristically generate high quality QAOA angles that are not necessarily globally optimal, given a specific set of $C(z)$ cost values for all feasible solutions. JuliQAOA has been shown to be a highly effective QAOA angle learning tool [13, 10, 44, 45]. One basin hopping iteration is used for each QAOA round, and $p$ is incremented in step sizes of $1$. Angle extrapolation is also used within each JuliQAOA $p$ angle optimization pass – where $p+1$ angles are initialized using the best angles found at the preceding $p$ step. Note that similar ideas have been used in previous studies to compute good QAOA angles at $p>1$ [46]. At each step of $p$ past $p=1$, the angle learning is repeated until the mean expectation value has improved over the previous step of $p$ – each additional angle learning step (e.g., at $p+1$) uses random angle initialization and since the learning is not necessarily optimal, different initializations can result in different learned QAOA angle quality. Similar sets of heuristic rules that has been used in JuliQAOA have been developed in other studies, motivated by similar ideas of iterative or recursive exploration of the $p+1$ energy landscape using the information gained from the $p$ energy landscape [47, 48, 46]. Importantly, this JuliQAOA simulation is performing full state vector simulation, meaning the training procedure has at least one advantage of not being hindered by shot noise (e.g., finite sampling effects). However, because of the size of the systems, in particular the $27$ qubit Ising models, this simulation is quite memory intensive, but also very compute time intensive, using both HPC resources and many weeks of elapsed wall clock time (including re-attempts at optimization where the expectation value does not improve monotonically). The aim of this study is to make use of this pre-computation of the QAOA angles at small system sizes, where such state-vector based training is at least tractable, and apply these parameters to much larger QAOA circuits.
@@
The heavy-hex graph $G=(V,E)$ is defined by a set of vertices $V$ and edges $E$. The Hamiltonian is defined by single-spin terms indexed by $i\in V$, the spin-spin interactions are indexed by the edges $(i,j)\in E$, and lastly the local cubic terms are indexed by $(i,j,k)$. The cost function input a vector of spins $z=(z_{1},z_{2},\ldots,z_{n})\in\{+1,-1\}^{n}$. Every heavy-hex lattice is bipartite and the bipartition can be defined as $V=V_{2}\sqcup V_{3}$ with $E\subset V_{2}\times V_{3}$, where $V_{3}$ consists of vertices of maximum degree $3$, and $V_{2}$ consists of vertices of maximum degree $2$. $W$ is the set of vertices in $V_{2}$ where each node has degree $2$, and the spins in this set $W$ define the center spin of each of the geometrically local cubic terms in Eq. (2), indexed as $(i,j,k)$. All variable states are spins, where the states are either $\{+1,-1\}$, and the polynomial coefficients, $d_{i},d^{(2)},d^{(3)}$ are randomly (and independently) chosen to be either +1+1 or 1-1, making these a type of hardware-compatible disordered $\pm J$ models. The QAOA circuits use a direct spin to qubit mapping; each qubit represents a single spin in the logical Ising model.
@@
Figure 15 in Appendix A gives a circuit diagram translation from the hardware-compatible Ising model to the corresponding QAOA circuit diagram, where the most expensive component of the circuit in terms of gate complexity is the phase separator. The QAOA circuit simulations run on the IBM hardware do not use any dynamical decoupling, and do not use any type of quantum error mitigation. The QAOA circuits rely on computing a 3-edge-coloring of the heavy-hex hardware graph – the 3-edge-coloring used in all of the hardware simulations is randomly generated using minimum edge coloring heuristic algorithms from Networkx [55]. That fixed edge coloring is then used in all $p$ layers (this is recalculated for each Ising model instance). In principle, these edge colorings could be randomized to generate circuit invariants with slightly different noise profiles when executed on the hardware, but we leave this for future study. Additionally, these edge colorings could be optimized to have distinct (average) gate-times in each layer, thus reducing the overall circuit time [56]. During compilation testing, we did attempt this type of 3-edge-coloring partitioning. However, in practice, the two qubit gate times of the IBM processors used in this study have very nearly uniform times, which means that no significant reduction of circuit execution time can be gained, and randomization of the two qubit 3-edge-coloring does not yield variance in total circuit durations. In principle, if a quantum computer had two qubit gate timings that were highly variable, an optimized staggered edge-coloring (for example, 3 colors which correspond to short, median, and long gate times) could be used to compile circuits such that the total circuit duration was minimized. In other words, if the two qubit gate timings effectively replicated a 3-edge-coloring via distinct timing bins, then compiling to that edge coloring could yield substantial improvements in total circuit duration.
@@
All of the quantum hardware simulations are performed using IBM Quantum programmable fixed-frequency superconducting transmon qubit [57] devices, which are summarized in Table 1. These devices have the heavy-hex two-qubit connectivity graph [49]. The quantum circuits are defined and executed using the Qiskit Python 3 library [58, 59]. The circuits are transpiled using a circuit optimization level of $3$ before being sent to the quantum computer. Qiskit versions 1.2.2, 0.45.0, and 2.1.2 were used across all of the circuit executions on the IBM QPUs. The circuits are run on the IBM quantum computer in the form of discrete jobs – for all devices the shot count per job is 1,000,0001{,}000{,}000, with the exception of ibm_torino where because of device usage limitations only $5,000$ shots were run per job (thus requiring more jobs to be run to get an equivalent total shot count to the other devices). One of the consequences of this is higher QPU compute time used in total for ibm_torino because there is a time cost associated with running many more individual jobs. Occasionally, for some of the IBM QPUs a job would fail to execute due to an unknown internal error – in these cases, the job was simply re-submitted to the backend (failed jobs are not counted towards reported aggregate statistics such as QPU time).
@@
For the Ising model where $C_{\text{Max}}$ is the global maximum possible energy and $C_{\text{Min}}$ is the global minimum possible energy. $C_{e}$ is the cost function $C(z)$, Eq. (2), also called the energy, of a single spin configuration. This approximation ratio formula is maximized for minimum energies, and minimized for maximum energies, meaning that this is intended to be applied to minimization optimization problems.
@@ II.1 Projected Entangled Pair States Simulation Methods
Appendix C compares different tensor network contraction methods, and we find that belief propagation (BP) offers a very good convergence of local expectation values, while being more efficient than alternative methods. Therefore, we use BP-based simple update [65, 66] (which is equivalent to the original simple update algorithm [67]) to simulate the time evolution of the PEPS representation of the quantum system, and BP to perform the final contraction of the tensor network for calculation of the expectation values or sampling (unless specifically stated otherwise). This is corroborated by Refs. [68, 69, 70, 71], who found similar properties of simple update and BP specifically for digital circuits defined on the heavy-hex connectivity graph.
@@ II.2 Matrix Product State Simulation Methods
There exist other types of approximate simulation methods which have been able to approximate larger scale quantum circuits, specifically in the case of heavy-hex connectivity circuits, namely LOWESA [76, 77] and MPS [68]. For this study, we focus on using PEPS simulations because PEPS is particularly suited for 2D system simulations. However, we do evaluate how well MPS and LOWESA perform on simulating these QAOA circuits for two purposes. First, at low circuit depth, we see very good agreement between all three methods, thus validating the accuracy of all three simulators in the low-entanglement/low-$p$ regime. But, at higher circuit depths, we see breakdown of numerical stability and convergence, in particular for LOWESA, thus highlighting the necessity of validating the QAOA simulations with PEPS. Appendix B details these LOWESA simulations.
@@ III Results
The results are organized into sub-sections as follows. First, in Section III.1 we present the trained QAOA angles, which we will then use for the remainder of the study as fixed QAOA parameters and evaluate how well they work for direct parameter transfer on much larger problem sizes. Next, in Section III.2 validates how well some of these fixed parameters work when transferred onto new (unseen) test-problem instances, using exact classical statevector simulations (up to $27$ qubits).
Section III.3 presents extensive large-scale IBM NISQ computer QAOA executions, using the ensemble of transferred QAOA angles. This section includes extensive QAOA sampling result analysis, including reporting very large QAOA volume measures.
Section III.4 presents extensive classical tensor network simulations, specifically PEPS, of the high depth QAOA circuits, validating how well the QAOA parameter transfer works on Ising model problem sizes of up to $156$ spins. This includes ground-state sampling rates, and convergence analysis of the PEPS simulations. Complementing the PEPS simulations, Section III.5 then validates PEPS simulation accuracy using MPS tensor network simulations.

### v2

@@ I Introduction
The Quantum Approximate Optimization Algorithm (QAOA) [1, 2, 3, 4, 5] is designed to sample good solutions of combinatorial optimization problems. QAOA is relatively NISQ (Noisy Intermediate-Scale Quantum) [6] computer friendly in that it is feasible to implement the algorithm at the scale of near term noisy quantum processors (short depth, small qubit counts). There have been a considerable number of studies on the properties of QAOA using both numerical simulations and current quantum processors, in particular in comparison to existing classical algorithms that solve combinatorial optimization problems [7, 8, 9, 10, 11, 12, 13, 14]. One of the central challenges with the algorithmic study of QAOA is scaling sampling analysis with respect to number of variables and QAOA depth. This scaling requires either somewhat small scale numerical simulations (limited to tens of qubits for exact statevector computations, or up to hundreds of qubits using tensor network methods or similar types of approximate numerical techniques) or quantum hardware simulations which are subject to a variety of sources of error. QAOA requires hyper-parameter selection that in turn generally requires significant compute time, especially if this process is performed on the quantum device. However, the QAOA parameter learning is distinct from many other types of variational quantum algorithms (VQAs) [15] because the number of trainable parameters does not scale, per layer, with the number of variables in the problem, meaning that parameter consistency among problems with varying problem sizes, or coefficient distributions, can be examined. It turns out that QAOA parameters do indeed exhibit parameter consistency for problem instances that share some similarity, such as similar connection structure. This property has been studied and utilized for efficient QAOA parameter finding in numerous studies  [^1] The QAOA parameter similarity property comes by several different terms, including parameter transfer, problem instance independent QAOA parameters, fixed angles, parameter concentration, parameter fixing, size invariant QAOA parameters, fixed control parameters, fixed point QAOA, and transfer learning [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Many of these types of parameter transfer studies are able to condense QAOA angle usage into fixed angle formulas, at least for average cases of certain problem types and typically at fairly low depth [36]. If QAOA can be used in a non-variational learning manner, then the prospects of using QAOA for meaningfully sized problem instances on large scale error-corrected quantum computers are good – the variational learning overhead of QAOA can be quite significant. Parameter setting heuristics, namely from parameter concentration, will enable this approach to use QAOA at much larger scale [37]. There do exist other non-variational QAOA approaches, such as using linear-ramp adiabatic schedules [38, 39], which at sufficiently large $p$ guarantees optimality.
@@
Here we show through extensive numerical simulation some of the favorable characteristics of parameter transfer, and some of the limitations. More precisely, we use QAOA parameters trained on small, statevector simulation-addressable problem instances to execute full scale QAOA runs on several IBM quantum processors. Finally, we validate the QAOA angle performance using extensive tensor network simulations, in particular PEPS (Projected Entangled Pair States). The QAOA circuits and optimization problems that we use in this study are carefully designed to be extremely hardware compatible to the heavy-hex graph and are therefore very NISQ-computer-friendly see refs [40, 41]. The sparsity of the variable connections in these optimization problem instances makes them an excellent test case for studying fundamental algorithmic engineering questions about QAOA. In this case, we are focusing on high QAOA depth parameter transfer.
@@
In this study, the high-depth scaling of QAOA parameters is characterized using an ensemble of learned QAOA angles on small (e.g., less than $30$ qubits) problem instances whose parameters are then transferred to be used on much larger problem instances (up to $156$ qubits). The central question that is examined is how far parameter transfer QAOA angles can be used, with respect to $p$. This is an important question because although parameter transfer seems to be incredibly successful for finding good performing QAOA angles, there are limits to direct parameter training for parameter transfer – namely, for the task of transferring parameters up to larger problem sizes, better angles cannot be learned past an approximation ratio of $1$, and higher QAOA depths are generally required as problem sizes increase. It is not clear how well parameter transfer can be applied to various combinatorial optimization problem instances in general, but it is crucial to understand where transferred QAOA angles fail to perform, or perhaps work well in certain cases. Remarkably one of the findings of this study is that the intensely optimized QAOA angles, using the classical numerical statevector simulation tool JuliQAOA, recover the smooth-schedule properties of discretized quantum annealing (adiabatic)-type protocols.
@@ II Methods
The QAOA angle finding in this study is carried out using the Julia [44] based software package JuliQAOA [45], where the angle finding is performed using basin hopping optimization, which in turn uses gradient descent, and all angle optimization begins with uniformly random guesses at $p=1$. JuliQAOA is an efficient numerical simulation angle finding tool that can heuristically generate high quality QAOA angles that are not necessarily globally optimal, given a specific set of $C(z)$ cost values for all feasible solutions. As far as we are aware, finding globally optimal QAOA parameters is a very hard computational problem; we are not aware of tools or techniques that would allow us to verify global optimality of QAOA parameters (e.g., up to some precision), even for relatively small optimization problem instance sizes. JuliQAOA has been shown to be a highly effective QAOA angle learning tool [13, 10, 46, 47]. One basin hopping iteration is used for each QAOA round, and $p$ is incremented in step sizes of $1$. Angle extrapolation is also used within each JuliQAOA $p$ angle optimization pass – where $p+1$ angles are initialized using the best angles found at the preceding $p$ step. Note that similar ideas have been used in previous studies to compute good QAOA angles at $p>1$ [48]. At each step of $p$ past $p=1$, the angle learning is repeated until the mean expectation value has improved over the previous step of $p$ – each additional angle learning step (e.g., at $p+1$) uses random angle initialization and since the learning is not necessarily optimal, different initializations can result in different learned QAOA angle quality. Similar sets of heuristic rules used in JuliQAOA have been developed in other studies, motivated by similar ideas of iterative or recursive exploration of the $p+1$ energy landscape using the information gained from the $p$ energy landscape [49, 50, 48, 51]. Importantly, this JuliQAOA simulation is performing full state vector simulation, meaning the training procedure has at least one advantage of not being hindered by shot noise (e.g., finite sampling effects). However, because of the size of the systems, in particular the $27$ qubit Ising models, this simulation is quite memory and compute time intensive. The angle training used both High Performance Computing (HPC) resources and many weeks of elapsed wall clock time (including re-attempts at optimization where the expectation value does not improve monotonically). The aim of this study is to make use of this pre-computation of the QAOA angles at small system sizes, where such state-vector based training is at least tractable, and apply these parameters to much larger QAOA circuits. We have found that this particular QAOA angle learning tool is the most effective tool that we are aware of; other optimization techniques, such as gradient-free optimization methods, fail systematically at high QAOA depths.
@@
The heavy-hex graph $G=(V,E)$ is defined by a set of vertices $V$ and edges $E$. The Hamiltonian is defined by single-spin terms indexed by $i\in V$, the spin-spin interactions are indexed by the edges $(i,j)\in E$, and lastly the local cubic terms are indexed by $(i,j,k)$. The cost function takes as input a vector of spins $z=(z_{1},z_{2},\ldots,z_{n})\in\{+1,-1\}^{n}$. Every heavy-hex lattice is bipartite and the bipartition can be defined as $V=V_{2}\sqcup V_{3}$ with $E\subset V_{2}\times V_{3}$, where $V_{3}$ consists of vertices of maximum degree $3$, and $V_{2}$ consists of vertices of maximum degree $2$. $W$ is the set of vertices in $V_{2}$ where each node has degree $2$, and the spins in this set $W$ define the center spin of each of the geometrically local cubic terms in Eq. (2), indexed as $(i,j,k)$. All variable states are spins, where the states are either $\{+1,-1\}$, and the polynomial coefficients, $d_{i},d^{(2)},d^{(3)}$ are randomly (and independently) chosen to be either +1+1 or 1-1, making these a type of hardware-compatible disordered $\pm J$ models. The complexity class, for exact optimal solution computation, that the Ising models of the form Eq. 2 is not known [57]. The QAOA circuits use a direct spin to qubit mapping; each qubit represents a single spin in the logical Ising model.
@@
Fig. 13 in Appendix A gives a circuit diagram translation from the hardware-compatible Ising model to the corresponding QAOA circuit diagram, where the most expensive component of the circuit in terms of gate complexity is the phase separator. The QAOA circuit simulations run on the IBM hardware do not use any dynamical decoupling, and do not use any type of quantum error mitigation. The QAOA circuits rely on computing a 3-edge-coloring of the heavy-hex hardware graph – the 3-edge-coloring used in all of the hardware simulations is randomly generated using minimum edge coloring heuristic algorithms from NetworkX [58]. That fixed edge coloring is then used in all $p$ layers (this is recalculated for each Ising model instance). In principle, these edge colorings could be randomized to generate circuit invariants with slightly different noise profiles when executed on the hardware, but we leave this for future study. Additionally, these edge colorings could be optimized to have distinct (average) gate-times in each layer, thus reducing the overall circuit time [59]. During compilation testing, we did attempt this type of 3-edge-coloring partitioning. However, in practice, the two qubit gate times of the IBM processors used in this study have very nearly uniform times, which means that no significant reduction of circuit execution time can be gained, and randomization of the two qubit 3-edge-coloring does not yield variance in total circuit durations. In principle, if a quantum computer had two qubit gate timings that were highly variable, an optimized staggered edge-coloring (for example, 3 colors which correspond to short, median, and long gate times) could be used to compile circuits such that the total circuit duration was minimized. In other words, if the two qubit gate timings effectively replicated a 3-edge-coloring via distinct timing bins, then compiling to that edge coloring could yield substantial improvements in total circuit duration.
@@
All of the quantum hardware simulations are performed using IBM Quantum programmable fixed-frequency superconducting transmon qubit [60] devices, which are summarized in Table 1. These devices have the heavy-hex two-qubit connectivity graph [52]. The quantum circuits are defined and executed using the Qiskit Python 3 library [61, 62]. The circuits are transpiled using a circuit optimization level of $3$ before being sent to the quantum computer. Qiskit versions 1.2.2, 0.45.0, and 2.1.2 were used across all of the circuit executions on the IBM QPUs. The circuits are run on the IBM quantum computer in the form of discrete jobs – for all devices the shot count per job is 1,000,0001{,}000{,}000, with the exception of ibm_torino where because of device usage limitations only 5,0005{,}000 shots were run per job (thus requiring more jobs to be run to get an equivalent total shot count to the other devices). One of the consequences of this is higher QPU compute time used in total for ibm_torino because there is a time cost associated with running many more individual jobs. Occasionally, for some of the IBM QPUs a job would fail to execute due to an unknown internal error – in these cases, the job was simply re-submitted to the backend (failed jobs are not counted towards reported aggregate statistics such as QPU time).
@@
For the Ising model where $C_{\text{Max}}$ is the global maximum possible energy and $C_{\text{Min}}$ is the global minimum possible energy. $C_{e}$ is the cost function $C(z)$, Eq. (2), also called the energy, of a single spin configuration. This approximation ratio formula is maximized for minimum energies, and minimized for maximum energies, meaning that this is intended to be applied to minimization problems.
@@ II.1 Projected Entangled Pair States Simulation Methods
Appendix D compares different tensor network contraction methods, and we find that belief propagation (BP) offers a very good convergence of local expectation values, while being more efficient than alternative methods. Therefore, we use BP-based simple update [68, 69] (which is equivalent to the original simple update algorithm [70]) to simulate the time evolution of the PEPS representation of the quantum system, and BP to perform the final contraction of the tensor network for calculation of the expectation values or sampling (unless specifically stated otherwise). This is corroborated by Refs. [71, 72, 73, 74], who found similar properties of simple update and BP specifically for digital circuits defined on the heavy-hex connectivity graph.
@@ II.2 Matrix Product State Simulation Methods
There exist other types of approximate simulation methods which have been able to approximate larger scale quantum circuits, specifically in the case of heavy-hex connectivity circuits, namely LOWESA [79, 80] and MPS [71]. For this study, we focus on using PEPS simulations because PEPS is particularly suited for 2D system simulations. However, we do evaluate how well MPS and LOWESA perform on simulating these QAOA circuits for two purposes. First, at low circuit depth, we see very good agreement between all three methods, thus validating the accuracy of all three simulators in the low-entanglement/low-$p$ regime. But, at higher circuit depths, we see breakdown of numerical stability and convergence, in particular for LOWESA, thus highlighting the necessity of validating the QAOA simulations with PEPS. Appendix C details these LOWESA simulations.
@@ III Results
The results are organized into sub-sections as follows. First, in Section III.1 we present the trained QAOA angles, which we will then use for the remainder of the study as fixed QAOA parameters and evaluate how well they work for direct parameter transfer on much larger problem sizes. Next, Section III.2 validates how well some of these fixed parameters work when transferred onto new (unseen) test-problem instances, using exact classical statevector simulations (up to $27$ qubits).
Section III.3 presents extensive large-scale IBM NISQ computer QAOA executions, using the ensemble of transferred QAOA angles. This section includes extensive QAOA sampling result analysis, including reporting large QAOA volume measures.
Section III.4 presents extensive classical tensor network simulations, specifically PEPS, of the high depth QAOA circuits, validating how well the QAOA parameter transfer works on Ising model problem sizes of up to $156$ spins. This includes ground-state sampling rates, and convergence analysis of the PEPS simulations. Complementing the PEPS simulations, Section III.5 then further validates PEPS simulation accuracy using MPS tensor network simulations. Combined, Section III.4 and Section III.5 (as well as Appendix C) show strong numerical evidence that large scale parameter transfer QAOA would work, up to high depth $p$ QAOA, on quantum computers with significantly lower error than the actual quantum computer hardware results in Section III.3.